OpenStudy (anonymous):

A baseball player hits a baseball into the air with an initial vertical velocity of 48 feet per second from a height of 3 feet. a) Write an equation that gives the baseball's height as a function of the time (in seconds) after it is hit. b) After how many seconds does the baseball reach its maximum height? c) What is the maximum height?

OpenStudy (anonymous):

\[y = v _{0} t \times \sin a - \frac{ g t ^{2} }{ 2 }\] maybe that helps

OpenStudy (anonymous):

a.) Distance in the vertical= vinitial*time+1/2(acceleration due to gravity*(time)^2) b.)Take derivative and set = 0, then solve for t. c.)plug value for t in a.) and solve for distance in the vertical.

OpenStudy (anonymous):

btw your only given a vertical velocity, so voila!

OpenStudy (anonymous):

accleration due to gravity (g) is a constant, 9.8m/s^2. The baseball acts opposite gravity therefore it has a negative acceleration.

OpenStudy (anonymous):

is this right for a)? \[h=-16t^{2}+48t+3\]

OpenStudy (jim766):

yeah I think so...

OpenStudy (jim766):

do you know how to find time till max hieight?

OpenStudy (anonymous):

no not really

OpenStudy (jim766):

H = Ax^2 + Bx + C is standard from A = -16 B = 48 C = 3 max will be at -B/2A

OpenStudy (anonymous):

so would 1.5 be right?

OpenStudy (jim766):

Thats good.... that is how many seconds before the ball hits max

OpenStudy (jim766):

Now to find the max height , substitute 1.5 for x in the equation

OpenStudy (anonymous):

So, 39 feet would be the max height?

OpenStudy (jim766):

perfect!

OpenStudy (anonymous):

Thank you so much... You are a life saver!

OpenStudy (jim766):

yw